A jet lands on a aircraft carrier at 744 mile (20 m/s).
(a)what is its acceleration if it stops in 6 sec.
(b)what is the displacement of the plane which it is stopping.
1
Expert's answer
2020-03-02T10:25:35-0500

Given:

"v_i=20\:\rm m/s;"

"v_f=0\:\rm m/s;"

"t=6\:\rm s."

(a) The acceleration


"a=\frac{v_f-v_i}{t}=\frac{0-20}{6}=-3.3\:\rm m/s^2."

(b) The displacement


"d=\frac{v_f+v_i}{2}t=\frac{0+20}{2}\times 6=60\:\rm m."